find the value of
5 X 10^-5
_____________
25 X 10^3


Answer by Earlsdon(6294)

Dividing numbers with exponents.
(5 X 10^-5) / (25 X 10^3) You can do the integer part first:
5/25 = 1/5
Then do the exponential part: 10^-5 / 10^3 Dividing exponents means you have to subtract the exponents: -5 -3 = -8 So 10^-5/10^3 = 10^-8
Generally, we like to see positive exponents in the answer, so 10^-8 = 1/10^8
The final answer is: (1/5)X10^-8 = 1/(5X10^8)
If you have to write the answer in "standard notation" (without exponents), then it would become:
(1/5)X10^-8 = 0.2X10^-8 = 0.000000002